OpenStudy (anonymous):

The sum of the reciprocals of two consecutive integers is -9/20. find the integers

OpenStudy (mertsj):

x = first integer x+1 = next consecutive integer. 1/x = reciprocal of the first integer 1/(x+1)= reciprocal of the next consecutive integer. Write the equation that says the sum of the reciprocals is -9/20

OpenStudy (anonymous):

I got -5 and -4?